Q-4 Jennings Corp. has 1,000,000 shares of S1 par value stock authorized, 200,000 shares issued, and 150,000 shares outstanding. On June 1, 2010, Jennings Board of Directors declared a 10% stock dividend at a time that the stock carried a market value of S30. Prepare the journal entry required to record the transaction described above and then post it to the related T-accounts: Compute the number of shares outstanding after the June 1 2010 stock dividend. Jennings Corp. announced a 100% stock dividend on June 1, 2011 Prepare the journal entry required to record the transaction described above and then post it to the related T-accounts: Compute the number of shares outstanding after the June 1, 2011 stock dividend. Jennings Corp. announced a 2 for 1 stock split on June 1, 2012 Prepare the journal entry required to record the transaction described above and then post it to the related T-accounts:
Answer #1
June 1, 2010 Retained earnings (150000*10%*$30) 450000
Stock dividends distributable 450000
Stock dividends distributable 450000
Common stock (15000*$1) 15000
Additional paid in capital-Common stock 435000
Number of shares outstanding after the
stock dividend = 150000+150000*10% = 165000 shares
June 1, 2011 Retained earnings (165000*100%*$1) 165000
Stock dividends distributable 165000
Stock dividends distributable 165000
Common stock 165000
Number of shares outstanding after the stock
dividend = 165000+165000 = 330000
June 1, 2012 For a stock split no regular journal entry is required. Only a Memorandum
entry is made in the General Journal and the General Ledger, as below:
"On June 1, 2012, a 2-for-1 stock split was declared for the common stockholders. This will result in the number of issued and outstanding common shares to increase from 330,000 shares to 660,000 shares. Further, the par value per share will decrease from $1.00 to $0.50."
Number of shares outstanding after the stock
split = 330000*2/1 = 660000
RETAINED EARNINGS
06/01/2010--Beginning Balance
06/01/2010 -Stock dividend 450000
06/01/2011--Stock dividend 165000
STOCK DIVIDENDS DISTRIBUTABLE
450000 06/01/2010--Retained Earnings
06/01/2010--Common stock 15000
06/01/2010--APIC--Common stock 435000
450000 450000
06/01/2011--Common stock 165000 165000 06/01/2011--Retained Earnings
165000 165000
0 06/01/2011--Ending balance
COMMON STOCK
0 150000 Beginning balance
0 15000 06/01/2010--Stock dividend
0 165000
165000 06/01/2011--Beginning balance
165000 06/01/2011--Stock dividend
0 330000
330000 06/01/2011--Ending balance
06-01-2012--"On June 1, 2012, a 2-for-1 stock split was declared for the common stockholders. This will result in the number of issued and outstanding common shares to increase from 330,000 shares to 660,000 shares. Further, the par value per share will decrease from $1.00 to $0.50."
APIC-COMMON STOCK
0 435000 06/01/2010--Stock dividends distributable
435000 435000
435000 06/01/2011--Ending balance